The following are ABO phenotypes from a population sample of one hundred persons. The number of individuals representing the four phenotypic classes are: type A, twenty; type B, forty-eight; type AB, sixteen; type O, sixteen. What are the three frequencies of the three different alleles (IA, IB and i)? What do you have to assume about this population to determine the allele frequencies from this data?
Answer #1

To determine the allele frequencies from this data, it will be needed to assume that the population follows Hardy - Weinberg principle.

Genotype of blood group A can be iAiA or iAi

Genotype of blood group B can be iBiB or iBi

Genotype of blood group AB is iAiB

Genotype of blood group O is ii.

According to Hardy-Weinberg principle, the genotypic frequencies with 3 alleles will be -

p2+ 2pq + q2 + 2pr + r2 + 2pr = 1

p- allele frequency of A, q - allele frequency of B, 2pq- allele frequency of AB, r2- allele frequency of O

N ( p2 + 2pr ) = the number of type A people in the population,

N ( q2 + 2qr ) = the number of type B people in the population,

N ( 2pq ) = the number of type AB people in the population, and

N ( r2 ) = the number of type O people in the population.

N ( r2 ) = 16 = 100 X r2

r2 = 16/100 = 0.16

r = 0.4

p2 + 2pr = 20/ 200 = 0.1

p2 + 2p (0.4) = 0.1

p2 + 0.8 p - 0.1 = 0

According to the quadratic equation solving -

p = 0.16

p + q + r = 1

0.16 + 0.4 + q = 1

q = 1- 0.56

q = 0.44